If f is monotonic on (a,b) then show that f is of bounded variation on(a,b)

1
Expert's answer
2022-02-15T17:52:21-0500

If "\ V(f,[a,b])< \infty" , then "f" is of bounded variation on "[a,b]" .

"V(f,[a,b])=\sup \big\{\sum\limits_{i=1}^n\big|f(x_i)-f(x_{i-1})\big|, \text{where }a=x_0<x_1<…<x_n=b\}"


1) "f" is increasing

"\sum\limits_{i=1}^n\big|f(x_i)-f(x_{i-1})\big|= \sum\limits_{i=1}^n\big(f(x_i)-f(x_{i-1})\big)=f(x_1)-f(x_0)+f(x_2)-f(x_1)+…+f(x_n)-f(x_{n-1})=f(x_n)-f(x_0)=f(b)-f(a)"


So, "V(f,[a,b])=f(b)-f(a)<\infty" .

Therefore "f" is of bounded variation on "[a,b]" .


2) "f" is decreasing "\sum\limits_{i=1}^n\big|f(x_i)-f(x_{i-1})\big|= \sum\limits_{i=1}^n\big(f(x_{i-1})-f(x_{i})\big)=f(x_0)-f(x_1)+f(x_1)-f(x_2)+…+f(x_{n-1})-f(x_{n})=f(x_0)-f(x_n)=f(a)-f(b)"


So, "V(f,[a,b])=f(a)-f(b)<\infty" .

Therefore "f" is of bounded variation on "[a,b]" .